OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (anonymous):

a solution to an equation means the number (or numbers) that will make the equation true. for example, the equation x+2=5. x=9 is not a solution because (9)+2=5 is NOT a true statement. but x=3 is a solution because (3)+2=5 is a true statement.

OpenStudy (anonymous):

thank you